OpenStudy (anonymous):

Prove. If p divides a and p divides a^2+b^2, then p divides b

OpenStudy (amistre64):

do we know yet that: a mod k + b mod k = (a+b) mod k ??

OpenStudy (anonymous):

No

OpenStudy (anonymous):

We are covering the gcd

OpenStudy (anonymous):

I wa string to say that if p divides a and p divides a^2+b^2 the p must divide their difference but that doesnt seem to work

OpenStudy (amistre64):

if not, then we simply have the definition of 'divides' to work with p|a means that for some integer n: a = pn p|(a^2+b^2) means that for some integer k: a^2+b^2 = pk since a=pn, a^2 = p^2n^2 therefore p^2n^2 + b^2 = pk b^2 = pk - p^2n^2 can you work it from there?

OpenStudy (amistre64):

since p divides the right side, p divides the left side should come into play with that

OpenStudy (amistre64):

otherwise we might be able to work the quadratic in b

OpenStudy (anonymous):

okay but i am not sure how that shows that p divides b if p divides b^2

OpenStudy (amistre64):

bb = p (k-pn^2) bb = p (k-an) since integers are a closed set then k-an is also just some integer, let it be t bb = pt can we conclude from there? or do we need to go a different route ?

OpenStudy (anonymous):

okay, that makes sense. thanks